George has 2 younger nieces, Emma and Maribel- Emma is he older of the 2 girls. He was thinking about their ages the other day. He realized these facts:
• in four years, he'll be 3 times as old as Maribel is right now
• in five years, his age will equal to the sum of his nieces' ages at that point
• three times Emma's age now, minute Maribel's present age, tells George's age last year
Assume everybody has whole-number ages- no decimals here!
How old is everybody right now?

Answer by stanbon(75887)

George has 2 younger nieces, Emma and Maribel-
Emma is the older of the 2 girls.
He was thinking about their ages the other day. He realized these facts:
• in four years, he'll be 3 times as old as Maribel is right now
g+4 = 3m
-----------------------
• in five years, his age will equal to the sum of his nieces' ages at that point
g+5 = e+5 + m+5
g - e - m = 5
---------------------
• three times Emma's age now, minus Maribel's present age, tells George's age last year
g-1 = 3e - m
g - 3e + m = 1
----------------------

Assume everybody has whole-number ages- no decimals here!
How old is everybody right now?
-----
Rearrange:
g + 0 - 3m = -4
g - e - m = 5
g - 3e + m = 1
--------------------
I used a matrix program on my TI-84 to get
g = 29 yrs
e = 13 yrs
m = 11 yrs
---------------
Cheers,
Stan H.
-------------